# Fabrication and characterization of optical filters from polymeric aerogels loaded with diamond scattering particles

### v2

@@ 5 Conclusion
Acknowledgments The authors gratefully acknowledge the National Aeronautics and Space Administration under grant numbers NNX14AB76A and NNH18ZDA001N-18-APRA18-0008, as well as the National Science Foundation Division of Astronomical Sciences for their support of this work under Grant Numbers 0959349, 1429236, 1636634, and 1654494. K.H acknowledges that the material is based upon work supported by NASA under award number 80GSFC21M0002. We also gratefully acknowledge support under the Goddard Internal Research and Development (IRAD) program.
@@
Supplemental document See Supplemental document for additional detail on the chemical properties of filter samples fabricated as part of this work.
Table 3: Chemical properties of filter samples fabricated as part of this work.

### v3

@@ 6 Backmatter
Funding This work was supported by the National Aeronautics and Space Administration under grant numbers NNX14AB76A and NNH18ZDA001N-18-APRA18-0008, as well as the National Science Foundation Division of Astronomical Sciences for their support of this work under Grant Numbers 0959349, 1429236, 1636634, and 1654494. K.H acknowledges that the material is based upon work supported by NASA under award number 80GSFC21M0002. We also gratefully acknowledge support under the Goddard Internal Research and Development (IRAD) program.
@@
Acknowledgments The authors gratefully acknowledge Aerogel Technologies, LLC for use of their facilities and support for sample fabrication during the initial stages of this work.
Portions of this work were presented at the SPIE conference on Astronomical Telescopes + Instrumentation in 2022, under papers titled “Characterization of aerogel scattering filters for astronomical telescopes” and “Novel infrared-blocking aerogel scattering filters and their applications in astrophysical and planetary science observations.”
The authors declare no conflicts of interest.
Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
[…]
Supplemental document See Supplement 1 for supporting content.
Table 3: *
Supplement 1. Chemical properties of filter samples fabricated as part of this work.